OpenStudy (anonymous):

You accidently throw your car keys horizontally at 8.0 m/s form a cliff 64m high. how far from the base of the cliff should you look for the keys Read more: You accidently throw your car keys horizontally at 8.0 m/s form a cliff 64m high. how far from the base of the cliff should you look for the keys?

OpenStudy (anonymous):

Bummer. Let's see. How long would it take for the keys to fall 64m if we simply dropped them off the cliff? The answer can be found in the following equation\[y(t) = - \left (1 \over 2 \right) g t^2\] If we solve this for \(t\), we can find the horizontal distance the keys travel from this equation\[x(t) = v_x t\]

OpenStudy (anonymous):

\[R = u \sqrt{\frac{ 2H }{ g }}\]